OpenStudy (anonymous):

One more thing, how will I get the derivative of cos2(theta)?

hartnn (hartnn):

know the chain rule ?

OpenStudy (abb0t):

\[\cos^2(\theta)\] correct? Follow what @hartnn said.

hartnn (hartnn):

i think its \(\cos 2 \theta\)

OpenStudy (anonymous):

no, it was actually cos2(theta)

OpenStudy (abb0t):

I think it might be more clear if you put the 2 inside the parenthesis, just for future reference, but either way, you would still use the chain rule.

hartnn (hartnn):

so do u know derivative of cos theta ?

OpenStudy (anonymous):

it's -sin theta

hartnn (hartnn):

so do u know how to use chain rule to get derivative of cos 2theta ?

hartnn (hartnn):

\(f(g(x))' = f'(g(x))\times g'(x)\) so, \((\cos 2\theta )'=(-\sin 2\theta )\times (2\theta)'=....?\)

OpenStudy (anonymous):

ahhh...it's basically the chain rule

OpenStudy (anonymous):

I though I would have to use some trig identity

hartnn (hartnn):

thats what i first asked :P so what u got ?

hartnn (hartnn):

if u got -2sin 2theta, then its correct

OpenStudy (anonymous):

thanks! I did :)